Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report Page - 20120817

Observers : C. Inserra (QUB) Flora Cellier-Holzem (LPHNE) 
Data analysts on duty : M. Fraser, D. Wright, A. Pastorello, S.Taubenberger

Time shared with Benetti Large Program (BLP) also over the second three nights rather than the breakdown of firewalled nights - this was due to RA pressue ranges of targets and having to work around the weather. Data reduced, as much as possible with the PESSTO pipeline

Total time spent BLP on 20120817 : 0h 
Total time overall in Aug 2012 : 16h (out of total of about 34hr clear) 
Total time lost on 20120817: 7h 

All times CST and approximate
10:20 the clouds are moving away, but the humidity is constant at 94%. Dome never open.

23.30 humidity dropped, still waiting.
00.15 Dome open, focus done.
00.20 SN2012ca gm13+griz (PESSTO followup)
       seeing~1.4
01.25 PSN J19371572-4218271 PESSTO classification
      young type IIP, best match with SN1999em +5d at z~0.019        
      seeing~1.5
01.40 SSS120810:231802-560926 PESSTO classification
      Blue continuum, possible broad lines, need higher S/N
      seeing~1.3
02.00 SSS120807:233315-330852 PESSTO classification
     really young type IIP, best match with SN2004et +1d at z~0.03 or SN2005cs +3d. However it's a type IIP close to maximum light.
     seeing ~1.2-1.3
02.25 PS1-12baa/PTF12gzk/LSQ12dwl gm13 +UBVRi PESSTO followup
     seeing~1.2-1.3 strong wind. Good S/N, from a quick extraction it does not seem evolved so much since the last obs (10/08)
03.00 Dome closed AGAIN. This time strong wind with peaks of 80km/h
05.45 No way, the wind is still over the limit to open the dome, the twilight is approaching (05.53) and we need at least 20 minutes under the limit to open again the dome.
     Finish for tonight.
Comments